[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Mon Jun 05, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Tue Jun 6 00:35:23 UTC 2017


TAFB Shift Log for Mon Jun 05 2017 20:35:11 GMT-0400 (EDT)

Shift: Swing Shift

Pacific Forecaster: J Lewitsky

Atlantic Forecaster: A Levine

Surface Analyst: E Rivera-Acevedo

Technical Issues Summary:

GFE/NCP issues: Craig switched NWPS over to Comp02 for NH2/Atlantic  
(NH1/Pacific currently inoperable). JL and ALevine did a 18z test run which  
successfully returned into GFE after 1hr37min of running. The output  
is "cold started" since there has not been a run since 06z on 06/04,  
although data after the first ~48 hours looks reasonable after the "cold  
start period". ALevine did a 21z run which successfully completed as well,  
but the RUNLEN was forced to only 72 hours (likely the phantom boundary  
condition files not all the way in for 168 hours).

N-AWIPS issues: M. Sardi called at 2240z to advise us to NOT reboot any  
NAWIPS workstations until further notice. A patch was done on HSU "ridges"  
NAWIPS box and the workstation did not properly come back up initially. The  
same MAY occur to TAFB workstations until the fix can be applied. CUMULUS  
clock is running 2-3 minutes behind actual time / other workstations.

Other issues: The thermostat in TAFB ops area was once again turned up all  
the way/maximum to 90 degrees.

Miscellaneous issues: The Surface desk PC tower is making a constant frog  
like sound. A NCO trouble ticket was opened on the issue. TAFB received  
multiple "robo-calls" from 3-6pm on every extension in ops.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170606/b1fb2f25/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list